OpenStudy (anonymous):

how do i solve this: -2= log(3x+5)

OpenStudy (alekos):

If a = Logx then x = 10^a

OpenStudy (alekos):

this is the definition of a logarithm

OpenStudy (alekos):

use this rule to solve

OpenStudy (anonymous):

so then would it be x=10^-2

OpenStudy (alekos):

close. it would be 3x+5 = 10^-2

OpenStudy (alekos):

now solve for x

OpenStudy (anonymous):

3x+5=0.01 3x=-4.99 x=-1.6633

OpenStudy (alekos):

beautiful! well done

OpenStudy (anonymous):

thank you :)

OpenStudy (alekos):

your welcome heidi